The perimeter of any square is 420 cm,then the area of this square will be
Answer by Cromlix(4381)

Perimeter = 2 * length + 2 * width(all same size)
Therefore one side will equal 420/4 = 105 cm
Area = length * width = 105 * 105 = 11025cm^2
Hope this helps.
:-)